### v2

# Beyond Minimax Optimality: A Subgame Perfect Gradient Method

Abstract
The study of unconstrained convex optimization has historically been concerned with worst-case a priori convergence rates. The development of the Optimized Gradient Method (OGM), due to Drori and Teboulle [12], Kim and Fessler [19], marked a major milestone in this study, as OGM achieves the optimal worst-case convergence rate among all gradient-span first-order methods. However, this notion of worst-case optimality is relatively coarse and allows OGM to have worst-case performance even on instances where stronger convergence guarantees are possible. For example, OGM is known to converge at its worst-case rate even on the toy example $Lx^{2}/2$, where exact convergence in just two steps is possible.
We introduce a notion of optimality which is stronger than minimax optimality that requires a method to give optimal dynamic guarantees that exploit any “non-adversarialness” in the first-order oracle’s reported information. We then give an algorithm which achieves this stronger optimality notion: the Subgame Perfect Gradient Method (SPGM). SPGM is a refinement of OGM whose update rules and convergence guarantees are dynamically computed in response to first-order information seen during the algorithm’s execution. From a game-theoretic viewpoint, OGM can be seen as one side of a Nash Equilibrium for the “minimization game” whereas SPGM can be seen as one side of a Subgame Perfect Equilibrium for the same game. We also show that SPGM can be implemented with minimal computational and storage overhead in each iteration and provide a Julia implementation.
1 Introduction
This paper considers black-box first-order methods for smooth convex minimization. Let $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ be an $L$-smooth convex function with a minimizer $x_{\star}\in{\mathbb{R}}^{d}$. We wish to construct algorithms that effectively minimize the normalized suboptimality:
$\frac{f(x_{N})-f(x_{\star})}{L\|x_{0}-x_{\star}\|^{2}/2},$
where $x_{N}$ is the final iterate of the method and $x_{0}$ is the starting iterate of the method.
Classically, the design of first-order methods has focused on algorithms with good (or in some cases, optimal) a priori worst-case guarantees [28]. In [12, 19], a momentum based first-order method known as the Optimized Gradient Method (OGM) was exhibited, and this method was shown in [9] to achieve the best possible worst-case performance on the class of $L$-smooth convex functions. This is often referred to as minimax optimality. To be precise, it was shown in [19] that for any $L$-smooth convex function $f$, OGM guarantees
$\frac{f(x_{N})-f(x_{\star})}{L\|x_{0}-x_{\star}\|^{2}/2}\leq\frac{1}{\tau^{\text{O}}_{N}},$
where we define $\tau^{\text{O}}_{N}$ in (OGM recurrence). [^4] We state the (OGM recurrence) in terms of a sequence $\tau^{\text{O}}_{0},\dots,\tau^{\text{O}}_{N}$. This differs from the original presentation in [19, Equation 6.13] that involves a sequence $\theta_{0},\dots,\theta_{N}$. Equivalence between the two recurrences follows from the change of variables: $\tau^{\text{O}}_{i}=2\theta_{i}^{2}$ for $i=0,\dots,N-1$ and $\tau^{\text{O}}_{N}=\theta_{N}^{2}$. It was later shown in [9] that as long as $d\geq N+2$, there exists a function $f_{\textup{hard}}$ so that for any gradient span first-order method,
$\frac{f_{\textup{hard}}(x_{N})-f_{\textup{hard}}(x_{\star})}{L\|x_{0}-x_{\star}\|^{2}/2}\geq\frac{1}{\tau^{\text{O}}_{N}}.$
Recall, a first-order method is a gradient span method if at each iteration, $x_{i}\in x_{0}+\operatorname{span}\{\nabla f(x_{0}),\dots,\nabla f(x_{i-1})\}$.
$x$$f(x)$$x_{0}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$
$x$$f(x)$$x_{0}$$x_{1}$$x_{2}$
Figure 1: Left: The first five iterates of OGM on $f(x)=Lx^{2}/2$ with $x_{0}=1$. OGM produces $x_{4}\approx 0.304$. Right: The first three iterates of SPGM on $f(x)=Lx^{2}/2$ with $x_{0}=1$. After seeing the history =${\cal H}=\left\{(x_{0},f_{0},g_{0}),(x_{1},f_{1},g_{1})\right\}$, SPGM determines that $x_{\star}=0$ for all $L$-smooth convex functions agreeing with ${\cal H}$.
While OGM achieves the optimal a priori worst-case guarantee, its guarantees may be overly conservative, and its execution may be suboptimal when faced with non-adversarial functions. As a simple example, consider the quadratic function $f(x)=\frac{L}{2}x^{2}$ with initial iterate $x_{0}=1$ (see Figure 1). It is known [20, Theorem 5.1] that OGM’s behavior on this function satisfies $f(x_{N})-f(x_{\star})=\frac{L\left\lVert x_{0}-x_{\star}\right\rVert^{2}}{2\tau^{\text{O}}_{N}}$. However, OGM’s choice of $x_{2},\dots,x_{N}$ is suboptimal, as after seeing just the function and gradient values at $x_{0}$ and $x_{1}$, one can deduce that $x_{\star}=0$ for any $L$-smooth convex function agreeing with this history. We see that OGM provides optimal guarantees on “hard instances”, but fails to improve its guarantees on “easy instances”.
In practical settings, the first-order oracle simply reports first-order information from a fixed function and may, at times, reveal useful information that would otherwise not be present in an adversarial function. Thus, our goal is to construct a method that exploits useful information in a dynamic way to achieve optimal convergence on both “hard instances” and “easy instances” in a formal sense.
In order to make this notion of optimality precise, we consider a game-theoretic framework for evaluating first-order methods. This will allow us to state a criterion for a first-order method that demands that it capitalizes on useful information. We then construct a method, the Subgame Perfect Gradient Method (SPGM), that achieves this stronger notion of optimality. Here, the terminology “subgame perfect” references the game-theoretic notion of a “Subgame Perfect Equilibrium”, which will be defined shortly. We present the game-theoretic framework next and then state precisely SPGM’s performance guarantees in Theorem 1.
1.1 A game-theoretic framework for analysis of first order methods
We consider a zero-sum game, which we refer to as the minimization game. This game is played between two players, Alice and Bob, where Alice represents the first-order method and Bob represents the first-order oracle. We will say that a set of triples $\{(x_{i},f_{i},g_{i})\}_{i\in{\cal I}}$ indexed by some index set ${\cal I}$ is interpolable if there exists an $L$-smooth convex function $f$ so that $f$ is $L$-smooth and convex, and $f(x_{i})=f_{i}$ and $\nabla f(x_{i})=g_{i}$ for all $i\in{\cal I}$.
The game proceeds in $N+1$ rounds, indexed by $n=0,\dots,N$. In round $n$, Alice chooses a point $x_{n}\in x_{0}+\operatorname{span}\left\{g_{0},\dots,g_{n-1}\right\}$ representing a query point, and Bob responds with a pair $(f_{n},g_{n})$ with $f_{n}\in{\mathbb{R}}$ and $g_{n}\in{\mathbb{R}}^{d}$ in such a way that there exists a function $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ so that $\{(x_{i},f_{i},g_{i})\}_{i=0}^{n}$ is interpolable. At the end of round $N$, Bob additionally specifies a triple $(x_{\star},f_{\star},g_{\star})\in{\mathbb{R}}^{d}\times{\mathbb{R}}\times{\mathbb{R}}^{d}$ with $g_{\star}=0$ so that $\{(x_{i},f_{i},g_{i})\}_{i\in[0,N]\cup\left\{\star\right\}}$ is interpolable. Alice’s payoff is the reciprocal of the normalized suboptimality:
$\frac{L\left\lVert x_{0}-x_{\star}\right\rVert^{2}/2}{f(x_{N})-f(x_{\star})}.$
The subgames where Alice has a decision to make can be fully specified by the first-order history =${\cal H}=\left\{(x_{i},f_{i},g_{i})\right\}_{i=0}^{n-1}$, which is guaranteed to be interpolable. We will denote by $\mathcal{A}$ the set of subgames where it is Alice’s turn to make a decision. A deterministic strategy for Alice in the minimization game is a mapping from each subgame to a query point $x_{n}\in x_{0}+\operatorname{span}(\left\{g_{0},\dots,g_{{n-1}}\right\})$.
Similarly, the subgames where Bob has a decision to make can be fully specified by the first-order history =${\cal H}=\left\{(x_{i},f_{i},g_{i})\right\}_{i=0}^{n-1}$ together with some $x_{n}\in x_{0}+\operatorname{span}(\left\{g_{0},\dots,g_{{n-1}}\right\})$. We will denote by $\mathcal{B}$ the set of subgames where Bob must make a decision. A deterministic strategy for Bob is a mapping from each subgame $({\cal H},x_{n})\in{\cal B}$ with $\left\lvert{\cal H}\right\rvert\leq N-1$ to some $(f_{n},g_{n})$ so that ${\cal H}\cup\{(x_{n},f_{n},g_{n})\}$ is interpolable, and a mapping for each subgame $({\cal H},x_{N})\in{\cal B}$ with $\left\lvert{\cal H}\right\rvert=N$ to $(f_{N},g_{N},\,x_{\star},f_{\star})$ so that ${\cal H}\cup\{(x_{N},f_{N},g_{N}),(x_{\star},f_{\star},g_{\star})\}$ is interpolable, where $g_{\star}\coloneqq 0$.
For a given pair of strategies $(A,B)$, the payoff of the strategies (from Alice’s perspective) is the final value of
$\frac{L\left\lVert x_{0}-x_{\star}\right\rVert^{2}/2}{f(x_{N})-f(x_{\star})},$
where the query points are chosen according to $A$ and the oracle responses are chosen according to $B$.
1.2 Equilibrium notions in the minimization game
We say that a pair of strategies $(A,B)$ is in Nash Equilibrium if neither player can increase their payoff by changing strategies. In this game, one strategy for Alice would be to choose query points consistent with OGM (see Algorithm 1). The OGM guarantee states
$f(x_{N})-f(x_{\star})\leq\frac{L}{2\tau^{\text{O}}_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}.$
That is, Alice is guaranteed a payoff of at least $\tau^{\text{O}}_{N}$ regardless of Bob’s play. Here, $\tau^{\text{O}}_{N}$ is the result of the following recurrence:
$\tau^{\text{O}}_{n}$$\coloneqq\begin{cases}2&\text{if }n=0\\ \tau^{\text{O}}_{n-1}+1+\sqrt{1+2\tau^{\text{O}}_{n-1}}&\text{if }n\in[1,N-1],\\ \tau^{\text{O}}_{N-1}+\frac{1+\sqrt{1+4\tau^{\text{O}}_{N-1}}}{2}&\text{if }n=N,\end{cases}$ (OGM recurrence)
and grows asymptotically like $\tau^{\text{O}}_{N}\approx\frac{N^{2}}{2}$.
One strategy for Bob would be to report the function and gradient values from the $L$-smooth function $f_{\textup{hard}}$. This strategy is available as long as $d\geq N+2$. By [9, Theorem 3], Bob can guarantee that
$f_{N}-f_{\star}\geq\frac{L}{2\tau^{\text{O}}_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}$
regardless of Alice’s play, i.e., Alice is guaranteed a payoff of at most $\tau^{\text{O}}_{N}$ regardless of Alice’s play. Thus, these two strategies form a Nash Equilibrium; neither player can increase their payoff by deviating their stategies.
While Nash Equilibrium is perhaps the most popular notion of equilibrium in a game, stronger notions exist. A pair of strategies $(A,B)$ is said to be in Subgame Perfect Equilibrium if for every subgame, the strategies defined by restricting $A$ and $B$ to that subgame are in Nash Equilibrium. This notion is stronger than that of Nash Equilibrium (see [32] for details on game theory), and the situation described in Figure 1 shows that OGM does not have this stronger property.
This notion is practically important when the function we are trying to optimize is not adversarially chosen. We may model the non-adversarial nature of typical functions as Bob playing this game “suboptimally”, i.e. playing a strategy that is not part of a Nash Equilibrium for the game. Thus, a strategy for Alice which is part of a Subgame Perfect Equilibrium can be regarded as an algorithm which capitalizes on suboptimal play to the largest extent possible, while still accounting for the possibility that Bob may begin to play adversarially in the future.
The following theorem essentially states SPGM is part of a Subgame Perfect Equilibrium.
Theorem 1.
For any $0<n\leq N$ and any set of first-order history =${\cal H}=\left\{(x_{i},f_{i},g_{i})\right\}_{i\in[0,n-1]}$ generated by SPGM, the output $x_{N}$ of running SPGM for $N-n$ additional iterations satisfies
$f(x_{N})-f(x_{\star})\leq\frac{L}{2\hat{\tau}_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2},$
where $\hat{\tau}_{N}$ is the result of (SPGM recurrence). Additionally, if $d\geq N+2$, then there exists an $L$-smooth convex function $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ agreeing with the history ${\cal H}$ such that the output $x_{N}$ from any method satisfying
$x_{i}\in x_{0}+\operatorname{span}\left(\left\{g_{0},\dots,g_{i-1}\right\}\right),\qquad\forall i\in[n,N]$
satisfies
$f_{N}-f_{\star}\geq\frac{L}{2\hat{\tau}_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}.$
Remark 1.
Technically, SPGM is not a “strategy” for Alice as SPGM does not assign an action to certain game states, namely those that cannot occur as a result of running SPGM on some first-order oracle. However, because we are primarily concerned with the design of algorithms, such subgames are not relevant to us, as they would not result from choices made by the algorithm. It can be shown via backward induction that there is a Subgame Perfect Equilibrium $(A,B)$ so that $A$ extends SPGM in the sense that $A$ plays the same action as SPGM in all cases where SPGM is defined, but this is beyond the scope of this paper.
An identical result holds for the setting where the history from only the $k$ most recent iterations is stored. In this setting, it is natural to augment the first-order history ${\cal H}$ with a minimal amount of additional information $\tau_{i}\in{\mathbb{R}}$ and $z_{i+1}\in{\mathbb{R}}^{d}$ related to intermediate deductions about the unknown function $f$ made in the $i$th iteration. See Sections 2.2 and 3 for the motivation and definition for $\tau_{i}$ and $z_{i+1}$.
Theorem 2.
Given any set of first-order history =${\cal H}=\left\{(x_{i},f_{i},g_{i},\tau_{i},z_{i+1})\right\}_{i\in[n-k,n-1]}$ generated by SPGM, the output $x_{N}$ of running SPGM for $N-n$ additional iterations satisfies
$f_{N}-f_{\star}\leq\frac{L}{2\tau_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2},$
where $\tau_{N}$ is the result of (SPGM recurrence). Additionally, if $d\geq N+2$, then there exists an $L$-smooth convex function $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ agreeing with the history ${\cal H}$ such that the output $x_{N}$ from any method satisfying
$x_{i}\in\operatorname{aff}\left(\left\{x_{0}\right\}\cup\left\{x_{i},z_{i+1}\right\}_{i\in[n-k,n-1]}\right)+\operatorname{span}\left(\left\{g_{n-k},\dots,g_{i-1}\right\}\right),\qquad\forall i\in[n,N]$ (1)
satisfies
$f_{N}-f_{\star}\geq\frac{L}{2\tau_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}.$
The computational overhead of limited-memory variant of SPGM can be shown to be $O(k^{3}+dk)$ time per iteration (see Remark 3).
SPGM can be derived constructively through the Performance Estimation Program (PEP) framework [12, 37]. Specifically, SPGM is an optimized version of OGM where the quantities $(f_{i},x_{i},g_{i})_{i\in[0,n-1]}$, which are treated as variables in the PEP framework before the start of the optimization procedure, become constants at round $n$. Surprisingly, the computation required to implement the SPGM amounts to solving a single low-dimensional convex optimization problem in each iteration. Thus, the SPGM can be implemented in a numerically stable manner with minimum additional overhead.
1.3 Related work
Smooth convex optimization is a classic problem with, by now, a lengthy history. We summarize some relevant literature both classic and modern.
Gradient descent, which dates back at least to Cauchy et al. [6], is perhaps the simplest algorithm for smooth convex optimization. In this setting, gradient descent with stepsize $1/L$, i.e., the algorithm $x_{n}=x_{n-1}-\frac{1}{L}g_{n-1}$ can be shown to achieve suboptimality $f_{N}-f_{\star}\leq\frac{L\left\lVert x_{0}-x_{\star}\right\rVert^{2}}{2N}$ [13]. In a seminal result of Nesterov [30], a method was constructed with the guarantee $f_{N}-f_{\star}\leq\frac{2L\left\lVert x_{0}-x_{\star}\right\rVert^{2}}{N^{2}}$. An almost matching lower bound was given by Nemirovsky and Yudin [28] showing that $f_{N}-f_{\star}\leq\frac{2L\left\lVert x_{0}-x_{\star}\right\rVert^{2}}{N^{2}}$ is optimal up to constants. This acceleration, from $N$ in the denominator to $N^{2}$ in the denominator, also generalizes to problems with simple constraints, minimizing the maximum of smooth convex functions, and the composite optimization framework (see [31]).
Since the early 2010s, there has been an explosion of work seeking optimal methods in large part due to the Performance Estimation Program (PEP) framework initiated by Drori and Teboulle [12], Taylor et al. [37]. There, it was shown that the problem of proving a convergence rate for a Fixed-Step First-Order Method (FSFOM) can be rephrased as solving a semidefinite program. Formally, a FSFOM is a first-order method where stepsizes and/or momentum terms are constants that are chosen offline independent of the first-order oracle’s output. This opened the door to the computer-assisted design of first-order methods [12, 16]. By now, a significant list of minimax optimal or conjectured minimax optimal first-order methods have been discovered. In the game-theoretic framework of Section 1.1, minimax optimal corresponds to the best possible payoff for Alice that she can guarantee when playing against an optimal adversary.
The Optimized Gradient Method for smooth convex optimization was first observed numerically in [12] and proved analytically in [19]. It was proved to be minimax optimal in [9]. Adaptive versions of this method have been designed using linesearching [10, 35] as well as restarting [22] techniques. An optimized first-order method for strongly convex minimization was derived in [39, 8, 33], culminating in the development of the minimax optimal Information-Theoretic Exact Method of [36]. Optimized methods able to be applied to various constrained settings or settings with proximal terms have been developed as well [38, 21, 18, 4, 17]. The minimax optimal constant stepsize gradient descent scheme for smooth convex or strongly convex optimization was first observed numerically in [12] and proved in [34, 24]. Conjectured minimax optimal long stepsize gradient descent schemes with partial acceleration were first observed numerically in [16] with upper bounds proved analytically in [1, 2, 14, 15, 40, 41].
In conjunction with all of the above algorithmic developments, a significant body of literature has now evolved for producing lower bounds [29, 28, 9, 11]. Our analysis in Section 4 relies heavily on the zero-chain constructions of [11].
We emphasize that all results above are a priori minimax optimal or conjectured minimax optimal. In contrast, the guarantees of SPGM are dynamic minimax optimal. Additionally, the SPGM is not a FSFOM as its stepsizes and momentum terms are chosen dynamically at runtime.
1.4 Outline
We provide some background information on PEP and OGM in Section 2. Section 3 introduces the SPGM formally and proves its convergence guarantee (Theorem 3). The proof of Theorem 3 is constructive in the sense that it shows how the update rules defining SPGM can be derived naturally from the PEP framework. Section 4 constructs matching worst-case functions agreeing with given history (see Propositions 1 and 2). Together, Theorems 3, 1 and 2 constitute a proof for Theorems 1 and 2. We conclude with preliminary numerical experiments in Section 5 comparing SPGM with gradient descent, OGM, and Limited-Memory BFGS [25].
2 Preliminaries
2.1 Smooth convex minimization and performance estimation
This section summarizes background material on the Performance Estimation Programming (PEP) methodology, initiated by Drori and Teboulle [12], Taylor et al. [37], in the setting of unconstrained smooth convex optimization.
Our goal is to find an optimal solution to the unconstrained minimization problem
$\min_{x\in{\mathbb{R}}^{d}}f(x),$
where $f$ is known to be convex and $L$-smooth, and is assumed to have a minimizer.
Recall, a convex function $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is said to be $L$ smooth if it is differentiable and for all $x,y\in{\mathbb{R}}^{d}$,
$\left\lVert\nabla f(x)-\nabla f(y)\right\rVert\leq L\left\lVert x-y\right\rVert.$
Instead of thinking about the function $f$ itself, it will be useful to think about the collection of first-order data that a first-order method encounters: $\left\{(x_{i},f_{i},g_{i})\right\}_{i\in{\cal I}}$. Here, ${\cal I}$ is some index set and we use the shorthand $f_{i}=f(x_{i})$ and $g_{i}=\nabla f(x_{i})$. The following lemma gives a characterization of when a set of first-order data is consistent (or interpolable) by an $L$-smooth convex function.
Lemma 1.
Given a set of first-order data $\left\{(x_{i},f_{i},g_{i})\right\}_{i\in{\cal I}}$, there exists a $L$-smooth convex function $f$ satisfying
$f(x_{i})=f_{i}\qquad\text{and}\qquad\nabla f(x_{i})=g_{i},\qquad\forall i\in{\cal I}$
if and only if the quantity
$Q_{i,j}\coloneqq f_{i}-f_{j}-\left\langle g_{j},x_{i}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{i}-g_{j}\right\rVert^{2}$
is nonnegative for all $i,j\in{\cal I}$.
2.2 Optimized Gradient Method (OGM)
The OGM algorithm can be described as follows:
Given $L$-smooth convex function $f$, initial iterate $x_{0}$, iteration budget $N$
  • Define $\tau^{\text{O}}_{0}=2$ and $z_{1}=x_{0}-2g_{0}$
  • For $n=1,\dots,N$
    • Reveal $g_{n-1}=\nabla f(x_{n-1})$
    • Set $\phi_{n}\coloneqq\tau^{\text{O}}_{n-1}$
    • Define
      $\psi_{n}\coloneqq\begin{cases}1+\sqrt{1+2\phi_{n}}&\text{if }n\leq N-1\\ \frac{1+\sqrt{1+4\phi_{n}}}{2}&\text{else}\end{cases},\qquad\qquad\tau^{\text{O}}_{n}\coloneqq\phi_{n}+\psi_{n},$
      $x_{n}\coloneqq\frac{\phi_{n}}{\tau^{\text{O}}_{n}}\left(x_{n-1}-\frac{1}{L}g_{n-1}\right)+\frac{\psi_{n}}{\tau^{\text{O}}_{n}}z_{n},\qquad\text{and}\qquad z_{n+1}\coloneqq z_{n}-\frac{\psi_{n}}{L}g_{n}.$
Algorithm 1 OGM
Note that at the time that $z_{n+1}=z_{n}-\frac{\psi_{n}}{L}g_{n}$ is defined, the quantity $g_{n}$ is still unknown. What we mean by this definition is that $z_{n+1}$ is a formal expression in $g_{n}$ that becomes a concrete vector in ${\mathbb{R}}^{d}$ upon learning $g_{n}$ at the beginning of iteration $n+1$.
The performance of OGM can be understood in terms of an inductive statement that is maintained throughout the algorithm. The elegant analysis below was first given by [35] and presented clearly in [13, Theorem 4.4] and [33]. We begin the induction by setting $\tau^{\text{O}}_{0}=2$ and partially defining $z_{1}=x_{0}-\frac{2}{L}g_{0}$. Then, the following expression is guaranteed to be nonnegative:
$H_{0}\coloneqq\tau^{\text{O}}_{0}\left(f_{\star}-f_{0}+\frac{1}{2L}\left\lVert g_{0}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{1}-x_{\star}\right\rVert^{2}\geq 0.$
Indeed, one can check that $H_{0}=2Q_{\star,0}$. Next, at iteration $n\in[1,N-1]$, we may assume inductively that
$H_{n-1}\coloneqq\tau^{\text{O}}_{n-1}\left(f_{\star}-f_{n-1}+\frac{1}{2}\left\lVert g_{n-1}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{n}-x_{\star}\right\rVert^{2}\geq 0.$
Then, the definitions of $\psi_{n},\phi_{n},\tau^{\text{O}}_{n},x_{n},z_{n+1}$ in OGM and direct algebraic manipulations yields that
$H_{n}$ $\coloneqq\tau^{\text{O}}_{n}\left(f_{\star}-f_{n}+\frac{1}{2L}\left\lVert g_{n}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{n+1}-x_{\star}\right\rVert^{2}$
$=H_{n-1}+\psi_{n}Q_{\star,n}+\phi_{n}Q_{n-1,n}.$
Thus, $H_{n}$ is nonnegative and we may continue the induction.
Finally, at iteration $N$, we make a slight change in the choice of $\psi_{N}$ so that
$\tau^{\text{O}}_{N}\left(f_{\star}-f_{N}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{N+1}-x_{\star}\right\rVert^{2}$
$\qquad=H_{N-1}+\psi_{N}Q_{\star,N}+\phi_{N}Q_{N-1,N}$
is nonnegative. Thus, the output $x_{N}$ of OGM is guaranteed to satisfy
$f_{N}-f_{\star}\leq\frac{1}{\tau^{\text{O}}_{N}}\left(\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{N+1}-x_{\star}\right\rVert^{2}\right)\leq\frac{L}{2\tau^{\text{O}}_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}.$
3 Subgame perfect strategies for the Minimization Game
This section introduces the Subgame Perfect Gradient Method (SPGM), which is a generalization of OGM to the setting where some amount of first-order information is known (or learned throughout the course of SPGM’s execution). Throughout this and the next section, we will assume that we have completed $n-1$ iterations of our algorithm, and have produced the following data:
${\cal H}$ =$=\left\{(x_{i},f_{i},g_{i},\tau_{i},z_{i+1})\right\}_{i=n-k}^{n-1}\text{ where}$
$f_{i}$ $=f(x_{i}),$
$g_{i}$ $=\nabla f(x_{i}),$
$\tau_{i}$ $>0$
We also assume that we have produced guarantees that the following expression is known to be nonnegative for all $i\in[n-k,n-1]$:
$H_{i}\coloneqq\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}-x_{\star}\right\rVert^{2}.$
We may assume that $z_{i+1}\neq x_{0}$ for any $i\in[n-k,n-1]$, as Corollary 1 shows that if this is not the case, then $f_{i}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\leq f_{\star}$ and that $x_{i}-\frac{1}{L}g_{i}\in\operatorname*{arg\,min}_{x}f(x)$.
Here, taking $k=n$ corresponds to the situation where all first-order history is stored.
Given this information, we deduce that the convex function $f$ that we are attempting to minimize must belong to the following subclass of $L$-smooth convex functions:
${\cal F}_{{\cal H}}\coloneqq\left\{f:\,\begin{array}[]{l}f\text{ is convex and }L\text{-smooth}\\ f(x_{i})=f_{i},\,\forall i\in[n-k,n-1]\\ \nabla f(x_{i})=g_{i},\,\forall i\in[n-k,n-1]\\ \tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}-x_{\star}\right\rVert^{2}\geq 0,\\ \hskip 200.0003pt\forall i\in[n-k,n-1]\end{array}\right\}.$
SPGM will behave similarly to OGM in that it will inductively maintain a nonnegative expression $H_{n}$ in each iteration. However, where OGM requires $H_{n}$ to be nonnegative for any $L$-smooth convex $f$ satisfying $H_{n-1}\geq 0$, SPGM only needs this expression to be nonnegative for $f$ in the subclass ${\cal F}_{\cal H}$. Thus, SPGM can pick the quantity $\phi_{n}$ more aggressively. In particular, we will choose $\phi_{n}$ using the subproblem defined in (4).
In order to define (4), we will need to define a number of auxiliary quantities for notational convenience. Let $Z$ be indexed by $[n-k,n-1]$ with $i$th column $z_{i+1}-x_{0}$ and let $G$ be indexed by $[n-k,n-1]$ with $i$th column $\frac{1}{L}g_{i}$. Define the vectors $\tau,v,h,q\in{\mathbb{R}}^{k}$ indexed by $[n-k,n-1]$:
$\tau_{i}$ $\coloneqq\tau_{i}$
$v_{i}$ $\coloneqq f_{i}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}$
$h_{i}$ $\coloneqq\tau_{i}v_{i}-\frac{L}{2}\left\lVert x_{0}\right\rVert^{2}+\frac{L}{2}\left\lVert z_{i+1}\right\rVert^{2}$
$q_{i}$ $\coloneqq f_{i}-\left\langle g_{i},x_{i}\right\rangle+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}.$
Finally, in terms of these quantities, we define $\phi_{n}$ to be the optimal value of the following nonnegative second-order cone problem in variables $\mu,\lambda_{\star}\in{\mathbb{R}}^{k}$ indexed by $[n-k,n-1]$:
$\phi_{n}$ $\coloneqq\sup_{\mu,\lambda_{\star}\in{\mathbb{R}}^{k}}\left\{\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle:\,\begin{array}[]{l}\frac{L}{2}\left\lVert Z\mu-G\lambda_{\star}\right\rVert^{2}\leq\left\langle\mu,h-v_{m}\tau-LZ^{\intercal}x_{0}\right\rangle+\left\langle\lambda_{\star},q-v_{m}1+LG^{\intercal}x_{0}\right\rangle\\ \mu,\lambda_{\star}\geq 0\end{array}\right\}.$ (4)
Lemma 2.
The problem (4) is feasible and its optimal value is at least $\tau_{n-1}$. If $z_{i+1}\neq x_{0}$ for any $i\in[n-k,n-1]$, then strong duality holds for (4).
Proof.
Consider the solution $\mu=\alpha e_{n-1}$ and $\lambda_{\star}=0$ for some $\alpha\in[0,1]$. We have that,
$\left\langle\mu,h-v_{m}\tau-Z^{\intercal}x_{0}\right\rangle+\left\langle\lambda_{\star},q-v_{m}1+G^{\intercal}x_{0}\right\rangle-\frac{L}{2}\left\lVert Z\mu-G\lambda_{\star}\right\rVert^{2}$
$\qquad=\alpha h_{n-1}-\alpha v_{m}\tau_{n-1}-L\alpha\left\langle z_{n}-x_{0},x_{0}\right\rangle-\frac{L\alpha^{2}}{2}\left\lVert z_{n}-x_{0}\right\rVert^{2}$
$\qquad=\alpha\tau_{n-1}(v_{n-1}-v_{m})+\frac{L(\alpha-\alpha^{2})}{2}\left\lVert z_{n}-x_{0}\right\rVert^{2}$
$\qquad\geq 0.$
Thus, this solution is feasible for any $\alpha\in[0,1]$. The feasible solution at $\alpha=1$ has objective value $\tau_{n-1}$.
If $z_{n}\neq x_{0}$, then for any $\alpha\in(0,1)$, the expression in the display above is strictly positive. Thus, (4) is essentially strictly feasible, in the sense of [5, Definition 1.4.2]. This guarantees strong duality holds for (4). ∎
Finally, we come to the definition of SPGM in terms of the optimization problem (4).
Given $L$-smooth convex function $f$, initial iterate $x_{0}$, iteration budget $N$
  • Define $\tau_{0}=2$ and $z_{1}=x_{0}-\frac{2}{L}g_{0}$
  • For $n=1,\dots,N$
    • Reveal $g_{n-1}=\nabla f(x_{n-1})$ and update first-order information set ${\cal H}\coloneqq\left\{(x_{i},f_{i},g_{i},\tau_{i},z_{i+1})\right\}_{i={n-k}}^{n-1}$
    • Let $m\in\operatorname*{arg\,min}_{i\in[n-k,n-1]}\left(f_{i}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)$
    • If $z_{i+1}=x_{0}$ for some $i\in[n-k,n-1]$ or if (4) is unbounded, break and output $x_{m}-\frac{1}{L}g_{m}$. Else, let $(\mu,\lambda_{\star})$ be an optimal solution to (4) and set $\phi_{n}=\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle$ and $z^{\prime}=x_{0}+Z\mu-G\lambda_{\star}$.
    • Define
      $\psi_{n}\coloneqq\begin{cases}1+\sqrt{1+2\phi_{n}}&\text{if }n\leq N-1\\ \frac{1+\sqrt{1+4\phi_{n}}}{2}&\text{else}\end{cases},\qquad\qquad\tau_{n}\coloneqq\phi_{n}+\psi_{n},$
      $x_{n}\coloneqq\frac{\phi_{n}}{\tau_{n}}\left(x_{m}-\frac{1}{L}g_{m}\right)+\frac{\psi_{n}}{\tau_{n}}z^{\prime},\qquad\text{and}\qquad z_{n+1}\coloneqq z^{\prime}-\frac{\psi_{n}}{L}g_{n}.$
Algorithm 2 SPGM
Remark 2.
An equivalent and more practical definition of SPGM would place the definition of $z_{n}$ immediately after revealing $g_{n-1}$. We adopt the form above to parallel our eventual proof strategies for SPGM’s convergence guarantees. We caution that any implementation of SPGM should carefully ensure that $z_{n}$ is defined using the $\psi_{n-1}$ and not $\psi_{n}$.
The following theorem states that SPGM correctly maintains an appropriate inductive hypothesis in each iteration as long as $(\mu,\lambda_{\star})$ is feasible.
Theorem 3.
Suppose $\mu,\lambda_{\star}\in{\mathbb{R}}^{k}_{+}$ are feasible in (4) and set $\phi_{n}=\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle$ and $z^{\prime}=x_{0}+Z\mu-G\lambda_{\star}$. Let $\tau_{n},\,z_{n+1},\,x_{n}$ be defined according to SPGM. If $n\leq N-1$, then
$\tau_{n}\left(f_{\star}-f_{n}+\frac{1}{2L}\left\lVert g_{n}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{n+1}-x_{\star}\right\rVert^{2}\geq 0$
for all $f\in{\cal F}_{\cal H}$. If $n=N$, then
$\tau_{N}(f_{\star}-f_{N})+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{N+1}-x_{\star}\right\rVert^{2}\geq 0$
for all $f\in{\cal F}_{\cal H}$.
We defer the proof of Theorem 3 to the next subsection.
Remark 3 (Computational and storage overhead of SPGM).
Running SPGM requires storing the vectors $\tau,v,h,q\in{\mathbb{R}}^{k}$, the matrices $Z,G\in{\mathbb{R}}^{d\times k}$, and the points $x_{0},x_{n-k},\dots,x_{n-1}\in{\mathbb{R}}^{d}$. It is also useful to dynamically maintain the inner product matrices $Z^{\intercal}Z$, $G^{\intercal}G$, $G^{\intercal}Z\in{\mathbb{R}}^{k\times k}$. This amounts to a total storage overhead of $O(dk)$ assuming that $k\ll d$. The computational overhead in each iteration involves updating the inner product matrices in $O(dk)$ time and solving the convex quadratic optimization problem (4) in $2k$ variables. Assuming that $k\ll d$, the computational cost of solving the low-dimensional optimization problem is dominated by $O(dk)$.
The following corollary to Theorem 3 shows that if SPGM terminates early, then SPGM correctly identifies a minimizer of $f$.
Corollary 1.
If $z_{i+1}=x_{0}$ for some $i\in[n-k,n-1]$ or if the program (4) is unbounded, then
$x_{m}-\frac{1}{L}g_{m}\in\operatorname*{arg\,min}_{x}f.$
Proof.
First, suppose $z_{i+1}=x_{0}$ for some $i\in[n-k,n-1]$. Then, $H_{i}\geq 0$ is equivalent to
$\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)\geq 0.$
Thus,
$f\left(x_{m}-\frac{g_{m}}{L}\right)\leq f_{m}-\frac{\left\lVert g_{m}\right\rVert^{2}}{2L}=v_{m}\leq v_{i}\leq f_{\star}.$
Suppose that (4) does not admit an optimal solution. Then, its domain must be unbounded and there exists a nonzero $(\mu,\lambda_{\star})$ such that $(\alpha\mu,\alpha\lambda_{\star})$ is feasible for all $\alpha\geq 0$.
Any such $(\mu,\lambda_{\star})$ must satisfy $Z\mu-G\lambda_{\star}=0$ as otherwise the quadratic constraint in $\alpha$ would be violated for all $\alpha$ large enough.
Now, let $\alpha\geq 0$ and consider the guarantee of Theorem 3 for $(\alpha\mu,\alpha\lambda_{\star})$ using the second definition of $\psi$. We have that $z^{\prime}=x_{0}+\alpha(Z\mu-G\lambda_{\star})=x_{0}$ and
$\phi_{n}$ $\coloneqq\alpha(\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle)$
$\psi_{n}$ $\coloneqq\frac{1+\sqrt{1+4\phi_{n}}}{2}$
$\tau_{n}$ $\coloneqq\phi_{n}+\psi_{n}$
$x_{n}$ $\coloneqq\frac{\phi_{n}}{\tau_{n}}\left(x_{m}-\frac{1}{L}g_{m}\right)+\frac{\psi_{n}}{\tau_{n}}x_{0}$
$z_{n+1}$ $\coloneqq x_{0}-\frac{\psi_{n}}{L}g_{n}.$
Fix $f\in{\cal F}_{{\cal H}}$. Then, by Theorem 3, for any $\alpha>0$,
$f(x_{n})-f_{\star}\leq\frac{1}{\tau_{n}}\left(\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert x_{0}-\psi_{n}g_{n}-x_{\star}\right\rVert^{2}\right)\leq\frac{L}{2\tau_{n}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}.$
Note that $\phi_{n}\to\infty$ as $\alpha\to\infty$. This follows as $(\mu,\lambda_{\star})$ is nonzero and $\tau$ and $1$ are entrywise positive. Thus, $x_{n}$ goes to $x_{m}-g_{m}/L$ as $\alpha\to\infty$ and the LHS converges to $f(x_{m}-g_{m}/L)-f_{\star}$. On the other hand, $\tau_{n}\to\infty$ as $\alpha\to\infty$. We conclude that $x_{m}-\frac{1}{L}g_{m}\in\operatorname*{arg\,min}f$. ∎
While Theorem 3 provides a guarantee
$f(x_{n})-f(x_{\star})\leq\frac{L}{2\tau_{n}}\|x_{0}-x_{\star}\|^{2}+\frac{1}{2}\|g_{n}\|^{2}$
on the quality of $x_{n}$ that can be made at iteration $n$, we will be more interested in a guarantee on the quality of $x_{N}$ that can be made at iteration $n$. To this end, we define the following sequence, which we can think of as being generated by the OGM update rule for $N-n$ iterations, where the starting bound is given by $\tau_{n}$ rather than $\tau_{0}$:
$\hat{\tau}_{n}$ $=\tau_{n}$ (SPGM recurrence)
$\hat{\tau}_{i}$$\coloneqq\begin{cases}\hat{\tau}_{i-1}+1+\sqrt{1+2\hat{\tau}_{i-1}}&\text{if }n<i<N\\ \hat{\tau}_{N-1}+\frac{1+\sqrt{1+4\hat{\tau}_{N-1}}}{2}&\text{if }n<i=N.\end{cases}$ (5)
The next corollary shows that $\hat{\tau}_{N}$ will be a lower bound on $\tau_{N}$ that is valid at iteration $n$ of SPGM.
Corollary 2.
Suppose that (4) has an optimal solution $(\mu,\lambda_{\star})$. Let $x_{N}$ denote the output of running SPGM for $N-n$ additional iterations. It holds that
$f(x_{N})-f(x_{\star})\leq\frac{1}{2\hat{\tau}_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}.$
Proof.
We may assume SPGM does not terminate early as otherwise SPGM would have found a minimizer of $f$. We will show inductively that $\tau_{i}\geq\hat{\tau}_{i}$ for all $i\in[n,N]$. By definition, $\tau_{i}=\hat{\tau}_{i}$ at iteration $i=n$. In all subsequent iterations $i\in[n+1,N]$, by Lemma 2, SPGM will take $\phi_{i}\geq\tau_{i-1}\geq\hat{\tau}_{i-1}$ and
$\tau_{i}$ =$=\begin{cases}\phi_{i}+1+\sqrt{1+2\phi_{i}}&\text{if }i<N\\ \phi_{i}+\frac{1+\sqrt{1+4\phi_{i}}}{2}&\text{if }i=N\end{cases}$
$\geq\begin{cases}\hat{\tau}_{i-1}+1+\sqrt{1+2\hat{\tau}_{i-1}}&\text{if }i<N\\ \hat{\tau}_{i-1}+\frac{1+\sqrt{1+4\hat{\tau}_{i-1}}}{2}&\text{if }i=N\end{cases}$
$=\hat{\tau}_{i}.$
The inequality follows as the expression in either case is an increasing function of $\phi_{i}$. ∎
3.1 Proof of Theorem 3
This subsection contains a proof Theorem 3.
Let $H_{n}$ denote the expression that we would like to certify is nonnegative:
$H_{n}$ $\coloneqq\tau_{n}\left(f_{\star}-f_{n}+\frac{1}{2L}\left\lVert g_{n}\right\rVert^{2}1_{n<N}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{n+1}-x_{\star}\right\rVert^{2}.$
Here, $1_{n<N}$ is the indicator function that takes the value $1$ if $n<N$ and is 00 otherwise.
The most direct proof simply expands the following claimed identity
$H_{n}$ $=\sum_{i=n-k}^{n-1}\mu_{i}H_{i}+\sum_{i=n-k}^{n-1}\lambda_{\star,i}Q_{\star,i}+\psi_{n}Q_{\star,n}+\phi_{n}Q_{m,n}+\epsilon,$
where $\epsilon\geq 0$ denotes the difference between the RHS and LHS in the first constraint in (4), and notes that the expressions on either side agree.
We now give a longer, but constructive proof that we hope clarifies the “magic” parameter choices in SPGM and makes future extensions of this analysis straightforward. By assumption, we have that for each $i\in[n-k,n-1]$,
$H_{i}=\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}-x_{\star}\right\rVert^{2}\geq 0.$
We now wish to select $\tau_{n}$, $x_{n}$ and $z_{n+1}$ such that for any choice of $g_{n}$, $f_{n}$, $x_{\star}$ and $f_{\star}$, the inequality $H_{n}\geq 0$ holds. Here, we assume that $z_{n+1}$ takes the form $z_{n+1}=z^{\prime}-\frac{\alpha}{L}g_{n}$ for some $\alpha$ and $z^{\prime}$ are independent of $g_{n},\,f_{n},\,x_{\star},\,f_{\star}$. Our strategy will be to write $H_{n}$ as a conic combination of the expressions $H_{i}$ for $i\in[n-k,n-1]$ and $Q_{i,j}$ for $i,j\in[n-k,n]\cup\left\{\star\right\}$, which are nonnegative by assumption, and the constant $1$. As the quantities $(x_{i},f_{i},g_{i},\tau_{i},z_{i+1})$ for $i\in[n-k,n-1]$ and $x_{0}$ are known, we can view the expressions $H_{i}$ and $Q_{i,j}$ as polynomials in the variables $g_{n}$, $f_{n}$, $x_{\star}$ and $f_{\star}$ only. Specifically, each expression takes the form
$C+af_{\star}+bf_{n}+\langle v,g_{n}\rangle+\langle w,x_{\star}\rangle+\frac{c}{2L}\|g_{n}\|^{2}+d\langle g_{n},x_{\star}\rangle+\frac{e}{2}\|x_{\star}\|^{2}$
for some choice of $C,a,b,c,d,e\in{\mathbb{R}}$ and $w,v\in{\mathbb{R}}^{d}$.
Our goal now becomes to find the $\tau_{n}$, $x_{n}$, $z^{\prime}$ and $\alpha$ so that the following equation holds as polynomials in these variables:
$H_{n}=\sum_{i=n-k}^{n-1}\mu_{i}H_{i}+\sum_{i=n-k}^{n}\lambda_{\star,i}Q_{\star,i}+\sum_{i=n-k}^{n-1}\lambda_{i,n}Q_{i,n}+\epsilon,$ (6)
where $\mu_{i},\lambda_{\star,i},\lambda_{i,n},\lambda_{\star,n},\,\epsilon\geq 0$ for each $i\in[n-k,n-1]$.
Remark 4.
Note that if $i,j\in[n-k,n-1]$, the quantity $Q_{i,j}$ is simply a nonnegative constant as it does not depend on any of the variables $g_{n},f_{n},x_{\star},f_{\star}$. For this reason, they can be safely omitted from our ansatz.
In principle, we could also include the inequalities $Q_{i,\star}\geq 0$ and $Q_{n,i}\geq 0$ in (6), which do involve $g_{n}$, $f_{n}$, $x_{\star}$ or $f_{\star}$. However, these inequalities “go the wrong way”, in the sense that they are either lower bounds on $f_{n}$ or upper bounds on $f_{\star}$. Moreover, our analysis later shows that in the worst case, these inequalities are slack.
We list the coefficient of each monomial in each of the expressions in (6) in Table 1.
$1$ $f_{\star}$ $f_{n}$ $g_{n}$ $x_{\star}$ $\frac{1}{2L}\|g_{n}\|^{2}$ $\langle g_{n},x_{\star}\rangle$
$H_{n}$ $\frac{L}{2}\|x_{0}\|^{2}-\frac{L}{2}\|z^{\prime}\|^{2}$ $\tau_{n}$ $-\tau_{n}$ $\alpha z^{\prime}$ $L(z^{\prime}-x_{0})$ $\tau_{n}1_{n<N}-\alpha^{2}$ $\alpha$
$H_{i}$ $-h_{i}$ $\tau_{i}$ $L(z_{i+1}-x_{0})$
$Q_{\star,i}$ $-q_{i}$ $1$ $-g_{i}$
$Q_{i,n}$ $v_{i}$ $-1$ $x_{n}-(x_{i}-\frac{1}{L}g_{i})$ $-1$
$Q_{\star,n}$ $1$ $-1$ $x_{n}$ $-1$ $1$
$1$ $1$
Table 1: Coefficients on each monomial in the expressions $H_{n},H_{i},Q_{\star,i},Q_{i,n},Q_{\star,n}$ for $i\in[n-k,n-1]$. Each row corresponds to an expression which makes up our certificate (where the index $i$ ranges from $n-k$ to $n-1$), and each column corresponds to a monomial. We have omitted the column $\|x_{\star}\|^{2}$ as its coefficient is zero in each of the expressions we consider.
We now equate the coefficients on each monomial on the LHS and RHS of (6) by taking a weighted sum across each column of Table 1. To ease notation, let $\lambda_{\star}\in{\mathbb{R}}^{k}$ denote the vector with entries $\lambda_{\star,n-k},\dots,\lambda_{\star,n-1}$. This produces the seven following equations:
$\frac{L}{2}\|x_{0}\|^{2}-\frac{L}{2}\|z^{\prime}\|^{2}$ $=-\left\langle h,\mu\right\rangle-\left\langle q,\lambda_{\star}\right\rangle+\sum_{i=n-k}^{n-1}\lambda_{i,n}v_{i}+\epsilon$ (7)
$\tau_{n}$ $=\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle+\lambda_{\star,n}$ (8)
$-\tau_{n}$ $=-\sum_{i=n-k}^{n-1}\lambda_{i,n}-\lambda_{\star,n}$ (9)
$\alpha z^{\prime}$ $=\sum_{i=n-k}^{n-1}\lambda_{i,n}\left(x_{n}-\left(x_{i}-\frac{1}{L}g_{i}\right)\right)+\lambda_{\star,n}x_{n}$ (10)
$L(z^{\prime}-x_{0})$ $=LZ\mu-LG\lambda_{\star}$ (11)
$\tau_{n}1_{n<N}-\alpha^{2}$ $=-\sum_{i=0}^{n-1}\lambda_{i,n}-\lambda_{\star,n}$ (12)
$\alpha$ $=\lambda_{\star,n}.$ (13)
Many of these inequalities simplify immediately. From equations (13), (11), (10), we deduce
$\alpha$ $=\lambda_{\star,n}$
$z^{\prime}$ $=x_{0}+Z\mu-G\lambda_{\star},\quad\text{and}$
$x_{n}$ $=\frac{1}{\lambda_{\star,n}+\sum_{i=n-k}^{n-1}\lambda_{i,n}}\left(\lambda_{\star,n}z^{\prime}+\sum_{i=n-k}^{n-1}\lambda_{i,n}\left(x_{i}-\frac{1}{L}g_{i}\right)\right).$
Now, we wish to maximize the value of
$\tau_{n}=\sum_{i=n-k}^{n-1}\lambda_{i,n}+\lambda_{\star,n}=\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle+\lambda_{\star,n}$
subject to the remaining constraints:
$\max_{\lambda_{i,n},\lambda_{\star},\lambda_{\star,n},\mu,\,\epsilon}\left\{\sum_{i=n-k}^{n-1}\lambda_{i,n}+\lambda_{\star,n}:\,\begin{array}[]{l}\sum_{i=n-k}^{n-1}\lambda_{i,n}=\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle\\ 1_{n<N}\left(\sum_{i=n-k}^{n-1}\lambda_{i,n}+\lambda_{\star,n}\right)-\lambda_{\star,n}^{2}=-\sum_{i=0}^{n-1}\lambda_{i,n}-\lambda_{\star,n}\\ \frac{L}{2}\|x_{0}\|^{2}-\frac{L}{2}\|x_{0}+Z\mu-G\lambda_{\star}\|^{2}=-\left\langle h,\mu\right\rangle-\left\langle q,\lambda_{\star}\right\rangle+\sum_{i=n-k}^{n-1}\lambda_{i,n}v_{i}+\epsilon\\ \lambda_{\star,n},\mu_{i},\lambda_{\star,i},\lambda_{i,n},\,\epsilon\geq 0\qquad\forall i\in[n-k,n-1]\end{array}\right\}.$
We make a few final simplifications. First, we may use the second constraint to solve for $\lambda_{\star,n}$:
$\lambda_{\star,n}=\begin{cases}1+\sqrt{1+2\sum_{i=n-k}^{n-1}\lambda_{i,n}}&\text{if }n<N,\\ \frac{1+\sqrt{1+4\sum_{i=n-k}^{n-1}\lambda_{i,n}}}{2}&\text{if }n=N.\end{cases}$
Second, noting that $\lambda_{\star,n}$ is an increasing function of $\sum_{i=n-k}^{n-1}\lambda_{i,n}$, it suffices to drop the term $\lambda_{\star,n}$ from the objective function. Third, we claim that it suffices to consider only feasible solutions where $\lambda_{i,n}=0$ for all $i\neq m$ (recall $m$ is chosen in $\operatorname*{arg\,min}_{i}v_{i}$). To see this, consider any feasible solution. We may modify this feasible solution by setting $\lambda_{m,n}\leftarrow\sum_{i}\lambda_{i,n}$ and $\lambda_{i,n}\leftarrow 0$ for all $i\neq m$. The effect is that $\sum_{i}\lambda_{i,n}$ remains constant and $\sum_{i}\lambda_{i,n}v_{i}$ is nonincreasing. We increase $\epsilon$ as necessary to maintain the third constraint. Fourth, we drop the variable $\epsilon\geq 0$ to get an inequality constraint.
Thus, we may define away $\lambda_{m,n}\coloneqq\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle$ and consider the problem
$\max_{\lambda_{\star},\mu}\left\{\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle:\,\begin{array}[]{l}\frac{L}{2}\|x_{0}\|^{2}-\frac{L}{2}\|x_{0}+Z\mu-G\lambda_{\star}\|^{2}\geq-\left\langle h,\mu\right\rangle-\left\langle q,\lambda_{\star}\right\rangle+v_{m}\left(\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle\right)\\ \lambda_{\star},\,\mu\geq 0\end{array}\right\},$
which is equivalent to (4).
We recognize $\lambda_{m,n}$ and $\lambda_{\star,n}$ as the quantities $\phi_{n}$ and $\psi_{n}$ in SPGM. This completes the proof of Theorem 3.
4 Dynamic Lower Bounds
In this section, we construct a worst-case function in ${\cal F}_{\cal H}$ showing that no gradient-span first-order method can outperform the guarantee of Corollary 2. Section 4.1 provides key definitions and constructs triples $\{(x_{i},f_{i},g_{i})\}_{i\in{\cal I}}$ where $\mathcal{I}\coloneqq[n-k,N]\cup\left\{\star\right\}$. We can then define the following hard function in $f_{{\cal H}}\in{\cal F}_{\cal H}$:
$f_{\cal H}(y)=\max_{\alpha\in\Delta_{\mathcal{I}}}\left\{\frac{L}{2}\left\lVert y-\sum_{i\in\mathcal{I}}\alpha_{i}\left(x_{i}-\frac{1}{L}g_{i}\right)\right\rVert^{2}+\sum_{i\in\mathcal{I}}\alpha_{i}\left(f_{i}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)\right\},$ (14)
where $\Delta_{\mathcal{I}}$ is the simplex in ${\mathbb{R}}^{{\cal I}}$. Section 4.2 then proves that (i) $f_{\cal H}$ lies in ${\cal F}_{\cal H}$ and interpolates each triple $(x_{i},f_{i},g_{i})$ (i.e., $f_{\cal H}(x_{i})=f_{i}$ and $\nabla f_{\cal H}(x_{i})=g_{i}$) and (ii) that no sequence of $N-n$ additional steps satisfying (1) can produce an objective gap less than $\frac{L\|x_{0}-x_{\star}\|^{2}}{2\tau_{N}}$. Both of these properties correspond to checking certain inequalities in terms of $\{(x_{i},f_{i},g_{i})\}_{i\in{\cal I}}$ established by [37, 11].
Remark 5.
The presentation in Sections 4.1 and 4.2 is mostly algebraic, but stems from the following guiding intuition. Our goal is to construct a function $f_{\cal H}\in{\cal F}_{\cal H}$ that is hard for all gradient span first-order methods. Here, “hard” means that no gradient span first-order method can outperform the bound in Corollary 2. Note that the guarantee of Corollary 2 matches the guarantee for the hybrid algorithm that takes a single SPGM step followed by $N-n-1$ OGM steps (see Section 2.2). Thus, as a first step, we construct a function so that the guarantee of Corollary 2 is tight for this hybrid algorithm. This requires us to specify the oracle’s strategy against this fixed strategy, i.e., the responses $(x_{i},f_{i},g_{i})$ for $i\in[n,N]\cup\left\{\star\right\}$. It should be expected that this construction will be stated in terms of the hybrid algorithm’s internal parameters $(\phi_{i},\psi_{i},\tau_{i},z_{i+1})$. The responses $(x_{i},f_{i},g_{i})$ are chosen to ensure that $Q_{i,j}\geq 0$ for all $i,j\in[n-k,N]\cup\left\{\star\right\}$ and $H_{i}\geq 0$ for all $i\in[n-k,n-1]$. Additionally, since the effective proof of Corollary 2 (see Theorems 3 and 2.2) uses the inequalities $Q_{m,n},Q_{n,n+1},\dots,Q_{N-1,N},Q_{\star,n},\dots,Q_{\star,N},H_{n},\dots,H_{N}\geq 0$, in order to match this guarantee, a hard instance must set each of these expressions to zero. The natural construction with these properties is presented in Section 4.1. We prove in Proposition 1 that these definitions satisfy the interpolating conditions and conclude that the convergence guarantee of Corollary 2 is tight for the hybrid algorithm on the function (14). Finally, Proposition 2 checks that this function is not only hard for the hybrid algorithm, but also for any gradient span first-order method (including SPGM).
4.1 Construction
It suffices to prove a lower bound in the case where (4) is bounded. To construct this worst-case function, we inspect the dual to (4). The following lemma states the dual of (4) and interprets complementary slackness between (4) and its dual. Its proof is standard and is deferred to Appendix B.
Lemma 3.
The dual program to (4) is
$\inf_{\xi,y}\left\{\frac{2\left\lVert y\right\rVert^{2}}{\xi L}+\frac{\xi L}{2}\left\lVert x_{0}\right\rVert^{2}+2\left\langle x_{0},y\right\rangle:\,\begin{array}[]{l}\xi(-h+v_{m}\tau)-2Z^{\intercal}y\geq\tau\\ \xi(-q+v_{m}1)+2G^{\intercal}y\geq 1\\ \xi>0.\end{array}\right\}$ (18)
Here, the $\inf$ can be replaced by a $\min$ as long as (4) is bounded. Strong duality holds between (4) and (18). If $(\mu,\lambda_{\star})$ and $(\xi,y)$ are optimal solutions to (4) and (18) respectively, then
$\frac{-2y}{\xi L}=x_{0}+Z\mu-G\lambda_{\star}=z^{\prime}.$
Define $f_{\star}$ so that $\frac{1}{\xi}=v_{m}-f_{\star}$. We rewrite dual feasibility in terms of $f_{\star}$ and $z^{\prime}$ as
$\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}\right\rVert^{2}+L\left\langle z_{i+1}-x_{0},z^{\prime}\right\rangle\geq 0\qquad\forall i\in[n-k,n-1]$ (19)
$f_{\star}-f_{i}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}-\left\langle g_{i},z^{\prime}-x_{i}\right\rangle\geq 0\qquad\forall i\in[n-k,n-1].$ (20)
For notational convenience, define $\Delta=\left\lVert z^{\prime}_{n}-x_{0}\right\rVert^{2}$. By strong duality, i.e., equating the primal and dual optimal values, we see that
$\phi$ $=\frac{2\left\lVert y\right\rVert^{2}}{\xi L}+\frac{\xi L}{2}\left\lVert x_{0}\right\rVert^{2}+2\left\langle x_{0},y\right\rangle$
$=\frac{\xi L}{2}\left\lVert\frac{2y}{\xi L}+x_{0}\right\rVert^{2}$
$=\frac{\xi L\Delta}{2}.$
Rearranging this identity shows that $f_{\star}=v_{m}-\frac{L\Delta}{2\phi}$.
For $i\in[n,N]$, let $\hat{\phi}_{i},\hat{\psi}_{i},\hat{\tau}_{i}$ denote the quantities that would be defined by one iteration of SPGM followed by $N-n-1$ iterations of OGM. Namely,
$\hat{\phi}_{n}$ $=\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle$
$\hat{\phi}_{i}$ $\coloneqq\tau_{i-1},\qquad\forall i\in[n+1,N]$
$\hat{\psi}_{i}$$\coloneqq\begin{cases}1+\sqrt{1+2\hat{\phi}_{i}}&\text{if }n\leq i\leq N-1\\ \frac{1+\sqrt{1+4\hat{\phi}_{i}}}{2}&\text{else}\end{cases}$
$\hat{\tau}_{i}$ $\coloneqq\hat{\phi}_{i}+\hat{\psi}_{i},\qquad\forall i\in[n,N].$
For the hard function $f_{{\cal H}}$, it will turn out that these quantities will agree with the values of $\phi_{i},\psi_{i},\tau_{i}$ that arise from running SPGM on $f_{{\cal H}}$ for the remaining $N-n-1$ iterations, and so we will abuse notation somewhat and let $\phi_{i}=\hat{\phi}_{i}$, $\psi_{i}=\hat{\psi}_{i}$, $\tau_{i}=\hat{\tau}_{i}$ for $i\in[n,N]$.
Inductively define $\eta_{n},\dots,\eta_{N}$ so that
$\eta_{i}$ $=\frac{1}{2\phi_{i}\psi_{i}}\left(1+\sum_{j=n}^{i-1}\psi_{j}^{2}\eta_{j}\right)\qquad\forall i\in[n,N].$
Note that, by definition, $\eta_{n}=\frac{1}{2\phi_{n}\psi_{n}}$. These choices of $\eta_{i}$ satisfy the following properties (proof in Appendix A).
Lemma 4.
It holds that $\eta_{i}>0$ for all $i\in[n,N]$,
$\frac{1}{2\tau_{i}(\psi_{i}-1)}\leq\eta_{i}\leq\frac{1}{2\phi_{n}\psi_{i}}\qquad\forall\,i\in[n,N],$
$\tau_{i}(\psi_{i}-1)\eta_{i}\leq\tau_{j}(\psi_{j}-1)\eta_{j}\qquad\forall\,i<j\in[n,N],\quad\text{and}$
$(\psi_{i}-1)\eta_{i}\geq\psi_{j}\eta_{j}\qquad\forall\,i<j\in[n,N].$
Let $g_{n},\dots,g_{N}$ denote mutually orthogonal vectors in $\left(\operatorname{range}(Z)+\operatorname{range}(G)\right)^{\perp}$, where $g_{i}$ has norm
$\left\lVert g_{i}\right\rVert^{2}=L^{2}\eta_{i}\Delta.$
Define
$z_{i+1}$ =$=\begin{cases}z^{\prime}_{n}-\frac{\psi_{n}}{L}g_{n}&\text{if }i=n\\ z_{i}-\frac{\psi_{i}}{L}g_{i}&\text{if }i\in[n+1,N]\end{cases}$
$x_{n}$ $=\frac{\phi_{n}}{\tau_{n}}\left(x_{m}-\frac{1}{L}g_{m}\right)+\frac{\psi_{n}}{\tau_{n}}z^{\prime}$
$x_{i}$ =$=\begin{cases}\frac{\phi_{n}}{\tau_{n}}(x_{m}-\frac{1}{L}g_{m})+\frac{\psi_{n}}{\tau_{n}}z^{\prime}&\text{if }i=n\\ \frac{\phi_{i}}{\tau_{i}}(x_{i-1}-\frac{1}{L}g_{i-1})+\frac{\psi_{i}}{\tau_{i}}z_{i}&\text{if }i\in[n+1,N]\end{cases}$
$f_{i}$ $=v_{m}-\frac{L\Delta}{2}\left(\frac{1}{\phi_{n}}-\left(2\psi_{i}-1\right)\eta_{i}\right)\qquad\forall i\in[n,N]$
$=f_{\star}+\frac{L\Delta}{2}\left(2\psi_{i}-1\right)\eta_{i}$
$x_{\star}$ $=z_{N+1}.$
4.2 Lower bound guarantees
Define $f_{\cal H}$ according to (14) using the set $\left\{(x_{i},f_{i},g_{i})\right\}_{i\in[n-k,N]\cup\left\{\star\right\}}$ constructed in Section 4.1. The following propositions establish desirable properties of $f_{\cal H}$.
Proposition 1.
Suppose $(\mu,\lambda_{\star})$ is an optimal solution to (4) and $\left(\frac{1}{v_{m}-f_{\star}},z^{\prime}\right)$ is an optimal solution to (18). Then $f_{\cal H}\in{\cal F}_{\cal H}$ is $L$-smooth, convex, and for all $i\in[n-k,N]\cup\left\{\star\right\}$, has
$f(x_{i})=f_{i},\qquad\nabla f(x_{i})=g_{i},$
and for all $i\in[n-k,n-1]$, has
$\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}-x_{\star}\right\rVert^{2}\geq 0.$
Proposition 2.
Suppose $(\mu,\lambda_{\star})$ is an optimal solution to (4) and $\left(\frac{1}{v_{m}-f_{\star}},z^{\prime}\right)$ is an optimal solution to (18). Then $f_{\cal H}$ possess the zero-chain property of [11], that is, for any $j\in[n,N-1]$,
$y\in\operatorname{aff}\left(\left\{x_{0}\right\}\cup\left\{x_{i},z_{i+1}\right\}_{i\in[n-k,n-1]}\right)+\operatorname{span}\left(\left\{g_{n-k},\dots,g_{j-1}\right\}\right)$
$\quad\implies\nabla f_{\cal H}(y)\in\operatorname{span}\left(\left\{x_{i}-x_{0},z_{i+1}-x_{0}\right\}_{i\in[n-k,n-1]}\cup\left\{g_{n-k},\dots,g_{j}\right\}\right)\ .$ (21)
These two propositions establish that any gradient-span first-order method playing the remaining iterations $n$ through $N$ against the hard instance $f_{\cal H}$ will have its $N$th iterate lie in the affine subspace $\operatorname{aff}\left(\left\{x_{n-k},\dots,x_{n-1},z_{n-k+1},\dots,z_{n}\right\}\right)+\operatorname{span}\left(\left\{g_{n-k},\dots,g_{N-1}\right\}\right)$. Since $g_{N}$ is orthogonal to this subspace, any such first-order method’s final iterate’s objective gap must be at least $f_{N}-f_{\star}=\frac{L}{2\tau_{N}}\left\lVert x_{0}-x_{\star}\right\rVert^{2}$. Since $f_{\cal H}\in{\cal F}_{\cal H}$, this establishes our matching lower bound.
The following lemmas will be useful in proving Proposition 1 and Proposition 2.
Lemma 5.
Suppose $j<i$ with $i\in[n,N]$. Then,
$\left\langle g_{j},x_{i}-x_{j}\right\rangle$ =$=\begin{cases}\frac{\tau_{j}}{\tau_{i}}(\psi_{j}-1)\eta_{j}L\Delta-\psi_{j}\eta_{j}L\Delta&\text{if }j\geq n\\ \frac{\phi_{n}}{\tau_{i}}\left\langle g_{j},x_{m}-g_{m}-z_{n}^{\prime}\right\rangle+\left\langle g_{j},z_{n}^{\prime}-x_{j}\right\rangle&\text{if }j<n\end{cases}$
Proof.
We first expand
$\left\langle g_{j},x_{i}-x_{j}\right\rangle$ $=\left\langle g_{j},x_{i}-z_{i}\right\rangle+\left\langle g_{j},z_{i}-x_{j}\right\rangle.$
Note that by definition, $z_{i}=z_{n}^{\prime}-\sum_{\ell=n}^{i-1}\frac{\psi_{\ell}}{L}g_{\ell}$, thus the second term is
$\left\langle g_{j},z_{i}-x_{j}\right\rangle$ =$=\begin{cases}-\psi_{j}\eta_{j}L\Delta&\text{if }j\geq n\\ \left\langle g_{j},z_{n}^{\prime}-x_{j}\right\rangle&\text{if }j<n\end{cases}.$
For the first term, we use the inductive definition of $x_{i}$, $z_{i}$, and $\phi_{i}$ to write
$\left\langle g_{j},x_{i}-z_{i}\right\rangle$ $=\frac{\tau_{i-1}}{\tau_{i}}\left\langle g_{j},x_{i-1}-\frac{1}{L}g_{i-1}-z_{i-1}+\frac{\psi_{i-1}}{L}g_{i-1}\right\rangle.$
Note that if $i-1\geq n$ and $i-1>j$, then $g_{j}$ and $g_{i-1}$ are orthogonal and we get that $\left\langle g_{j},x_{i}-z_{i}\right\rangle=\frac{\tau_{i-1}}{\tau_{i}}\left\langle g_{j},x_{i-1}-z_{i-1}\right\rangle$. Thus, we can unroll this recursion to get
$\left\langle g_{j},x_{i}-z_{i}\right\rangle$ =$=\begin{cases}\frac{\tau_{j}}{\tau_{i}}(\psi_{j}-1)\eta_{j}L\Delta&\text{if }j\geq n\\ \frac{\phi_{n}}{\tau_{i}}\left\langle g_{j},x_{m}-\frac{1}{L}g_{m}-z_{n}^{\prime}\right\rangle&\text{if }j<n\end{cases}.\qed$
Proof of Proposition 1.
It suffices to check that $H_{i}\geq 0$ for all $i\in[n-k,n-1]$ and $Q_{i,j}\geq 0$ for all $i,j\in[n-k,N]\cup\left\{\star\right\}$. This breaks into the cases below:
Suppose $i\in[n-k,n-1]$, then
$H_{i}$ $=\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}-x_{\star}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}-x_{\star}\right\rVert^{2}$
$=\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}\right\rVert^{2}+L\left\langle z_{i+1}-x_{0},x_{\star}\right\rangle$
$=\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}\right)+\frac{L}{2}\left\lVert x_{0}\right\rVert^{2}-\frac{L}{2}\left\lVert z_{i+1}\right\rVert^{2}+L\left\langle z_{i+1}-x_{0},z^{\prime}_{n}\right\rangle$
$\geq 0.$
Here, the third line follows as $x_{\star}=z^{\prime}_{n}-\sum_{\ell=n}^{N}\frac{\psi_{\ell}}{L}g_{\ell}$ where $g_{\ell}$ is orthogonal to $z_{i+1}-x_{0}$ for all $\ell\in[n,N]$. The last line follows by (19).
If $i,j\in[n-k,n-1]$, then $Q_{i,j}$ contains only quantities given by ${\cal H}$. We have by assumption that $Q_{i,j}\geq 0$.
Suppose $i\in[n-k,n-1]$ and $j\in[n,N]$. Then,
$Q_{i,j}$ $=f_{i}-f_{j}-\left\langle g_{j},x_{i}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{i}-g_{j}\right\rVert^{2}$
$=f_{i}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}-f_{j}-\frac{\eta_{j}L\Delta}{2}$
$\geq v_{m}-f_{j}-\frac{\eta_{j}L\Delta}{2}$
$=\frac{L\Delta}{2}\left(\frac{1}{\phi_{n}}-\left(2\psi_{j}-1\right)\eta_{j}\right)-\frac{\eta_{j}L\Delta}{2}$
$=\frac{L\Delta}{2}\left(\frac{1}{\phi_{n}}-2\psi_{j}\eta_{j}\right)$
$\geq 0.$
Here, the first line follows as $g_{j}$ is orthogonal to $g_{i}$ and $x_{i}-x_{j}$. The second line follows as $m\in\operatorname*{arg\,min}_{i\in[n-k,n-1]}v_{i}$. The last inequality follows by Lemma 4.
Suppose $i\in[n-k,n-1]$ and $j=\star$. Then,
$Q_{i,\star}$ $=f_{i}-f_{\star}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}$
$\geq v_{m}-f_{\star}$
$>0.$
Here, the last line follows by dual feasibility.
Suppose $i\in[n,N]$ and $j\in[n-k,n-1]$. Then,
$Q_{i,j}$ $=f_{i}-f_{j}-\left\langle g_{j},x_{i}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{i}-g_{j}\right\rVert^{2}$
$=v_{m}-(v_{m}-f_{i})-f_{j}-\frac{\phi_{n}}{\tau_{i}}\left\langle g_{j},x_{m}-\frac{1}{L}g_{m}-z^{\prime}_{n}\right\rangle-\left\langle g_{j},z_{n}^{\prime}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}$
$=\frac{\tau_{i}-\phi_{n}}{\tau_{i}}\left(v_{m}-f_{j}-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}+\left\langle g_{j},x_{j}-z^{\prime}_{n}\right\rangle\right)-(v_{m}-f_{i})-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}$
$\qquad+\frac{\phi_{n}}{\tau_{i}}\left(v_{m}-f_{j}-\left\langle g_{j},x_{m}-x_{j}-\frac{1}{L}g_{m}\right\rangle-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}\right)$
$=\frac{\tau_{i}-\phi_{n}}{\tau_{i}}\left(f_{\star}-f_{j}-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}+\left\langle g_{j},x_{j}-z^{\prime}_{n}\right\rangle\right)+\frac{\phi_{n}}{\tau_{i}}Q_{m,j}$
$\qquad+\frac{\tau_{i}-\phi_{n}}{\tau_{i}}\left(v_{m}-f_{\star}\right)-(v_{m}-f_{i})-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}$
$\geq\frac{\tau_{i}-\phi_{n}}{\tau_{i}}\frac{L\Delta}{2\phi_{n}}-\frac{L\Delta}{2}\left(\frac{1}{\phi_{n}}-\left(2\psi_{i}-1\right)\eta_{i}\right)-\frac{\eta_{i}L\Delta}{2}$
$=\frac{L\Delta}{2}\left(2\left(\psi_{i}-1\right)\eta_{i}-\frac{1}{\tau_{i}}\right)$
$\geq 0.$
Here, the second line uses the fact that $g_{i}$ is orthogonal to $g_{j}$ and the definition of $f_{i}$ and Lemma 5. The first inequality recognizes that $f_{\star}-f_{j}-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}+\left\langle g_{j},x_{j}-z^{\prime}_{n}\right\rangle\geq 0$ by (20) and $Q_{m,j}\geq 0$ by assumption. The last line follows from Lemma 4.
Suppose $i\in[n,N]$ and $j=\star$. Then,
$Q_{i,\star}$ $=f_{i}-f_{\star}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}$
$=\frac{L\Delta}{2}\left((2\psi_{i}-1)\eta_{i}-\eta_{i}\right)$
$=L\Delta(\psi_{i}-1)\eta_{i}$
$\geq 0.$
Suppose $i=\star$ and $j\in[n-k,n-1]$. Then,
$Q_{\star,j}$ $=f_{\star}-f_{j}-\left\langle g_{j},x_{\star}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}$
$=f_{\star}-f_{j}-\left\langle g_{j},z^{\prime}_{n}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}$
$\geq 0.$
Here, the second line follows as $x_{\star}=z_{n}^{\prime}-\sum_{\ell=n}^{N}\frac{\psi_{\ell}}{L}g_{\ell}$ where $g_{\ell}$ is orthogonal to $g_{j}$ for all $\ell\in[n,N]$. The inequality follows from (20).
Suppose $i=\star$ and $j\in[n,N]$. Then,
$Q_{\star,j}$ $=f_{\star}-f_{j}-\left\langle g_{j},x_{\star}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}$
$=f_{\star}-f_{j}+\frac{\psi_{j}}{L}\left\lVert g_{j}\right\rVert^{2}-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}$
$=-(2\psi_{j}-1)\frac{\eta_{j}L\Delta}{2}+(2\psi_{j}-1)\frac{\eta_{j}L\Delta}{2}$
$=0.$
Suppose $i,j\in[n,N]$ with $i<j$. Then,
$Q_{i,j}$ $=f_{i}-f_{j}-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}$
$=\frac{L\Delta}{2}\left((2\psi_{i}-2)\eta_{i}-(2\psi_{j})\eta_{j}\right)$
$\geq 0.$
Here, the last line follows from Lemma 4.
Suppose $i,j\in[n,N]$ with $i>j$. Then,
$Q_{i,j}$ $=f_{i}-f_{j}-\left\langle g_{j},x_{i}-x_{j}\right\rangle-\frac{1}{2L}\left\lVert g_{i}\right\rVert^{2}-\frac{1}{2L}\left\lVert g_{j}\right\rVert^{2}$
$=\frac{L\Delta}{2}((2\psi_{i}-1)\eta_{i}-(2\psi_{j}-1)\eta_{j})-\frac{\tau_{j}}{\tau_{i}}(\psi_{j}-1)\eta_{j}L\Delta+\psi_{j}\eta_{j}L\Delta-\frac{\eta_{i}L\Delta}{2}-\frac{\eta_{j}L\Delta}{2}$
$=\frac{L\Delta}{\tau_{i}}\left(\tau_{i}(\psi_{i}-1)\eta_{i}-\tau_{j}(\psi_{j}-1)\eta_{j}\right)$
$\geq 0.$
Here, the second line follows from Lemma 5 and the last line follows from Lemma 4. ∎
Proof of Proposition 2.
Conditions ensuring that (14) satisfies (21) for $j\in[n,N-1]$ are given by Drori and Taylor [11, Theorem 3]. In particular, it suffices to show
$\langle g_{i},g_{j}\rangle=\langle g_{i},g_{\ell}\rangle,\qquad\forall j\in[n,N-1],i\in[0,j-1],\ell\in[j+1,N]$
$f_{j}-\frac{1}{L}\langle g_{i},g_{j}\rangle+\frac{1}{2L}\|g_{j}-L(x_{j}-x_{0})\|^{2}-\frac{L}{2}\|x_{j}-x_{0}\|^{2}$
$\quad\geq f_{\ell}-\frac{1}{L}\langle g_{i},g_{\ell}\rangle+\frac{1}{2L}\|g_{\ell}-L(x_{\ell}-x_{0})\|^{2}-\frac{L}{2}\|x_{\ell}-x_{0}\|^{2},\qquad\forall i\in[0,N],\forall n\leq j<\ell\leq N$
$g_{j}\text{ is linearly separable from }\{g_{j+1},\dots,g_{N}\},\qquad\forall j\in[n,N-1]$
The first and third conditions above hold immediately as the considered gradients are all nonzero and orthogonal to each other. Hence we only need to verify the second condition, which can be algebraically simplified to
$Q_{i,\ell}-Q_{i,j}+\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle\geq 0.$
We verify this below in several cases:
First suppose $i\in[0,n-1]$ and consider any $n\leq j<\ell\leq N$. Since $\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=0$, we have
$Q_{i,\ell}-Q_{i,j}+\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=f_{j}-f_{\ell}+\frac{(\eta_{j}-\eta_{\ell})L\Delta}{2}=L\Delta(\psi_{j}\eta_{j}-\psi_{\ell}\eta_{\ell})\geq 0$
which is nonnegative as $\psi_{j}\eta_{j}\geq\psi_{\ell}\eta_{\ell}$.
Next suppose $i\in[n,N]$. Considering any $i<j<\ell\leq N$, since $\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=0$, we have
$Q_{i,\ell}-Q_{i,j}+\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=L\Delta(\psi_{j}\eta_{j}-\psi_{\ell}\eta_{\ell})\geq 0$
where the inequality uses that $\psi_{j}\eta_{j}\geq\psi_{\ell}\eta_{\ell}$ by Lemma 4. Considering $i=j<\ell\leq N$, since $\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=0$,
$Q_{i,\ell}-Q_{i,j}+\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=Q_{i,\ell}-0+0\geq 0.$
Considering any $n\leq j<i<\ell\leq N$, since $\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=-\langle g_{j},x_{i}-x_{j}\rangle$, Lemma 5 ensures
$Q_{i,\ell}-Q_{i,j}+\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle$
$=\frac{L\Delta}{2}((2\psi_{i}-2)\eta_{i}-2\psi_{\ell}\eta_{\ell})-\frac{L\Delta}{\tau_{i}}(\tau_{i}(\psi_{i}-1)\eta_{i}-\tau_{j}(\psi_{j}-1)\eta_{j})-(\frac{\tau_{j}}{\tau_{i}}(\psi_{j}-1)\eta_{j}L\Delta-\psi_{j}\eta_{j}L\Delta)$
$=\Delta(\psi_{j}\eta_{j}-\psi_{\ell}\eta_{\ell})\geq 0.$
Considering any $n\leq j<i=\ell\leq N$, since $\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=-\langle g_{j},x_{i}-x_{j}\rangle$, Lemma 5 ensures
$Q_{i,\ell}-Q_{i,j}+\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle$
$=0-\frac{L\Delta}{\tau_{i}}(\tau_{i}(\psi_{i}-1)\eta_{i}-\tau_{j}(\psi_{j}-1)\eta_{j})-(\frac{\tau_{j}}{\tau_{i}}(\psi_{j}-1)\eta_{j}L\Delta-\psi_{j}\eta_{j}L\Delta)$
$=\Delta((\psi_{i}-1)\eta_{i}+\psi_{j}\eta_{j})\geq 0$
where the inequality follows as both component terms are nonnegative. Finally, consider any $n\leq j<\ell<i\leq N$. Since $\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle=\langle g_{\ell},x_{i}-x_{\ell}\rangle-\langle g_{j},x_{i}-x_{j}\rangle$, applying Lemma 5, we have
$Q_{i,\ell}-Q_{i,j}+\langle g_{\ell}-g_{j},x_{i}-x_{0}\rangle$
$=\frac{L\Delta}{\tau_{i}}(\tau_{i}(\psi_{i}-1)\eta_{i}-\tau_{\ell}(\psi_{\ell}-1)\eta_{\ell})-\frac{L\Delta}{\tau_{i}}(\tau_{i}(\psi_{i}-1)\eta_{i}-\tau_{j}(\psi_{j}-1)\eta_{j})$
$\qquad+(\frac{\tau_{\ell}}{\tau_{i}}(\psi_{\ell}-1)\eta_{\ell}L\Delta-\psi_{\ell}\eta_{\ell}L\Delta)-(\frac{\tau_{j}}{\tau_{i}}(\psi_{j}-1)\eta_{j}L\Delta-\psi_{j}\eta_{j}L\Delta)$
$=L\Delta(\psi_{j}\eta_{j}-\psi_{\ell}\eta_{\ell})\geq 0.\qed$
5 Numerical results
This section evaluates the computational performance of SPGM on unconstrained smooth minimization problems involving random and real data. We implement SPGM in Julia and use MOSEK [3] via JuMP [26] to solve (4) in each iteration of SPGM. All experiments are run on an Intel i7 processor with 64GB of memory. A GitHub repository containing our implementations and preliminary numerical experiments is available at
https://github.com/ootks/SubgamePerfectGradientMethod
We compare the performance of SPGM and its limited memory variant with the following benchmarks: gradient descent (GD) with constant stepsize, i.e., $x_{n}=x_{n-1}-\frac{1}{L}\nabla f(x_{n-1})$, the Optimized Gradient Method (OGM) [19], and BFGS and its limited-memory variant (L-BFGS) [25]. We implement GD and OGM directly in Julia and use the implementations of BFGS and L-BFGS (restricted to a memory of ten past iterations) provided in Optim.jl [27].
Overall, we find SPGM with either a full memory or a memory limited to size $k=10$ (denoted SPGM-10) consistently outperforms GD and OGM and their well-established worst-case smooth convex optimization guarantees by exploiting the non-adversarialness of the given problem instances. Moreover, SPGM and its associated guarantee $1/\tau_{n}$ stay fairly competitive with the convergence of BFGS and L-BFGS. Figure 2 shows the scaled objective gap $(f(x_{n})-f(x_{\star}))/\frac{L}{2}\|x_{0}-x_{\star}\|^{2}$ and $1/\tau_{n}$ for each method on two representative problem instances from the larger experiments below with random and real problem data respectively.
Figure 2: Two representative plots of convergence of $(f(x_{n})-f(x_{\star}))/\frac{L}{2}\|x_{0}-x_{\star}\|^{2}$ for the considered methods. Left: A randomly generated instance of minimizing the logSumExp function (26) with dimensions $d=256,m=1024$. Right: An instance of the logistic regression (28) using the LIBSVM dataset “ionosphere” with $d=34,m=351$. On these instances, the limited memory versions of SPGM and BFGS had nearly identical convergence to their full memory counterparts and are omitted.
The optimal subgame payoff, $\hat{\tau}_{N}$, of the remaining game after $n$ iterations is a function of $n$ that increases with each suboptimal or non-advarsarial response from the first-order oracle. Figure 3 plots $1/\hat{\tau}_{N}$ for SPGM and SPGM-10, as a function of $n$. We also plot the constant non-adaptive guarantee of OGM, $1/\tau^{\text{O}}_{N}$, for reference.
Figure 3: Two representative plots of $1/\hat{\tau}_{N}$ as a function of $n$ as SPGM-10 and SPGM are run on Least Squares Regression problems of dimension $d=512,m=2048$. OGM displays its constant guarantee, $1/\tau^{\text{O}}_{N}$, providing the baseline, non-adaptive guarantee. Left instance is of the form (22) with $N=300$ and right instance is of the form (25) with $N=30$.
5.1 A sample of randomly generated problem instances
We consider six different families of smooth convex minimization problems all parameterized by $A\in\mathbb{R}^{m\times d}$, $b\in\mathbb{R}^{m}$, and $x_{0}\in\mathbb{R}^{d}$. For each problem family, we draw seven problem instances $A,b,x_{0}$ with sampled with i.i.d. normal entries, each with fixed, different problem dimensions $d=8,16,32,\dots,512$ and $m=4d$. Together these provide a collection of $42$ synthetic problem instances to compare performance across. The problem instances considered fall within two general categories.
First, we consider (regularized) least squares regression problems defined in the following four ways:
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{m}\|Ax-b\|^{2}_{2},$ (22)
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{m}\|Ax-b\|^{2}_{2}+\frac{1}{2}\|x\|^{2}_{2},$ (23)
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{m}\|Ax-b\|^{2}_{2}+h_{100}(\|x\|_{2}),$ (24)
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{m}\|Ax-b\|^{2}_{2}+\sum_{i=1}^{n}h_{100}(|x_{i}|)$ (25)
where $h_{L}(r)$ takes the value $\frac{L}{2}r^{2}$ if $r\leq 1$ and the value $Lr-\frac{L}{2}$ otherwise, denotes the Huber function. Note $h_{100}(\|x\|_{2})$ provides a smooth approximation of $100\|x\|_{2}$ and $\sum_{i=1}^{n}h_{100}(|x_{i}|)$ provides a smooth approximation of regularization by $100\|x\|_{1}$. While the first two definitions have constant Hessian, the latter may not have Hessians exist everywhere. Second, we consider smoothed approximations of minimizing a finite maximum $\max_{i}\{a_{i}^{T}x-b_{i}\}$, defined in the following two ways:
$\min_{x\in\mathbb{R}^{n}}f(x)=\log\left(\sum_{i=1}^{m}\exp(a_{i}^{T}x-b_{i})\right),$ (26)
$\min_{x\in\mathbb{R}^{n}}f(x)=\rho_{\max}(Ax-b)$ (27)
where $\rho_{\max}(z)=\min_{z^{\prime}}\{\max_{i}z^{\prime}_{i}+\frac{1}{2}\|z^{\prime}-z\|^{2}_{2}\}$ denotes the Moreau envelope of the max function. Note while the first formulation is analytic, the second is not even twice differentiable everywhere.
Sampling seven instances from these six classes provides our first test set. Figure 4 shows the performance of each considered method, displaying at each iteration $n$, the faction of these $42$ problem instances for which the method has reached accuracy
[…]
corresponding to the method reaching a low, medium, or high level of (appropriately rescaled) accuracy.
[…]
5.2 Regression problem instances derived from real data
To provide a more realistic numerical survey, we next consider performance on two types of regression problems with data from the LIBSVM dataset [7]. We consider six regularized least squares regressions of the form (25) with data from the “housing”, “mpg”, “pyrim”, “space_ga”, “triazines”, “bodyfat” datasets. Further, we consider six regularized logistic regression problems defined as
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{m}\sum_{i=1}^{m}\log(1+\exp(b_{i}\cdot a_{i}^{T}x))+\frac{1}{2m}\|x\|^{2}_{2},$ (28)
using normalized data $A,b$ from the classification datasets “australian”, “diabetes”, “heart”, “ionosphere”, “splice”, and “sonar”. The performance on these twelve instances is presented in Figure 5.
Figure 5: Performance comparison over twelve instances derived from LIBSVM data for the regression problems (25) and (28). From left to right, the target accuracy ranges as $10^{-3},10^{-6},10^{-9}$. Often, the performance of SPGM-10 matched that of SPGM with full memory, and hence, the lines overlap.
6 Conclusion and discussion
This paper offers a framework for analyzing first-order methods beyond the classical notion of minimax optimality. In contrast to minimax optimality, which can be thought of as a strategy in a Nash Equilibrium for the minimization “game”, the notion of a Subgame Perfect Equilibrium requires that a first-order method optimally capitalize on any non-adversarial first-order information. We believe this is a fundamental avenue towards deriving formal beyond-worst-case results in the design of optimization algorithms.
As a first algorithm in this space, we design the Subgame Perfect Gradient Method (SPGM) for unconstrained smooth convex minimization. SPGM matches the performance guarantee of the Optimized Gradient Method when faced with an adversarial function, but improves upon these guarantees dynamically as first-order information is revealed. We show that SPGM’s dynamic convergence guarantees are optimal among algorithms that can adapt to past information. We conjecture that subgame perfect versions of known minimax optimal first-order methods such as ITEM for smooth strongly convex optimization [11] and OGM-G for minimizing the gradient norm [23] also exist and leave these extensions as future work.
Appendix A Proof of Lemma 4
Proof of Lemma 4.
We will require the following identity
$\tau_{i}(\psi_{i}-1)$ =$=\begin{cases}\phi_{i}\psi_{i}+\tau_{i}&\text{if }i\in[n,N-1],\\ \phi_{N}\psi_{N}&\text{if }i=N,\end{cases}$ (29)
which follows from $\tau_{i}=\psi_{i}^{2}/2$ for $i\in[n,N-1]$ and $\tau_{N}=\psi_{N}^{2}$. Additionally, for all $i\in[n,N-1]$, it holds that
$2\phi_{i+1}\psi_{i+1}\eta_{i+1}$ $=1+\sum_{\ell=n}^{i}\psi_{\ell}^{2}\eta_{\ell}$
$=1+\sum_{\ell=n}^{i-1}\psi_{\ell}^{2}\eta_{\ell}+\psi_{i}^{2}\eta_{i}$
$=2\phi_{i}\psi_{i}\eta_{i}+\psi_{i}^{2}\eta_{i}$
$=2\tau_{i}(\psi_{i}-1)\eta_{i}.$
Rearranging this identity and noting that $\phi_{i+1}=\tau_{i}$, we have that for all $i\in[n,N-1]$, that $\psi_{i+1}\eta_{i+1}=(\psi_{i}-1)\eta_{i}$.
We prove the first statement inductively. For $i=n$, we have that
$\frac{1}{2\tau_{n}(\psi_{n}-1)}$ $=\frac{1}{2(\phi_{n}\psi_{n}+\tau_{n})}\leq\eta_{n}=\frac{1}{2\phi_{n}\psi_{n}}.$
Now, suppose $i\in[n+1,N]$, then
$\frac{1}{2\tau_{i}(\psi_{i}-1)}$ $\leq\frac{1}{2\phi_{i}\psi_{i}}$
$\leq\eta_{i}=\frac{(\psi_{i-1}-1)\eta_{i-1}}{\psi_{i}}$
$\leq\frac{(\psi_{i-1}-1)}{2\phi_{n}\psi_{i-1}\psi_{i}}$
$\leq\frac{1}{2\phi_{n}\psi_{i}}.$
Here, the first line follows from (29).
For the second claim, suppose $i\in[n,N-2]$. Then,
$\tau_{i}(\psi_{i}-1)\eta_{i}$ $=\tau_{i}\psi_{i+1}\eta_{i+1}$
$=\phi_{i+1}\psi_{i+1}\eta_{i+1}$
$\leq\tau_{i+1}(\psi_{i+1}-1)\eta_{i+1},$
where the inequality follows from (29).
For the third claim, first suppose $i\in[n,N-1]$ and $j=i+1$. Then, $\psi_{j}\eta_{j}=(\psi_{i}-1)\eta_{i}$ so that the claim holds. Now, suppose $n\leq i<j\leq N$. Then, we can chain together this inequality to get
$\psi_{j}\eta_{j}$ $\leq(\psi_{j-1}-1)\eta_{j-1}$
$\leq\psi_{j-1}\eta_{j-1}$
$\leq(\psi_{j-2}-1)\eta_{j-2}$
[…]
$\leq(\psi_{i}-1)\eta_{i}.\qed$
Appendix B Proof of Lemma 3
Proof of Lemma 3.
The primal problem (4) can be written as
$\sup_{\mu,\lambda_{\star}\in{\mathbb{R}}^{k}}\left\{\left\langle\tau,\mu\right\rangle+\left\langle 1,\lambda_{\star}\right\rangle:\,\begin{array}[]{l}\begin{pmatrix}\left\langle\mu,h-v_{m}\tau-LZ^{\intercal}x_{0}\right\rangle+\left\langle\lambda_{\star},q-v_{m}1+LG^{\intercal}x_{0}\right\rangle&Z\mu-G\lambda_{\star}\\ Z\mu-G\lambda_{\star}&\frac{2}{L}I\end{pmatrix}\succeq 0\\ \mu,\lambda_{\star}\geq 0\end{array}\right\}.$
Define dual variables
$\begin{pmatrix}\xi&(\frac{L\xi}{2}x_{0}+y)^{\intercal}\\ \frac{L\xi}{2}x_{0}+y&\Xi\end{pmatrix}\succeq 0,\qquad a\geq 0,\qquad b\geq 0.$
A direct computation of the dual gives
$\inf_{\xi,y,\Xi,a,b}\left\{\frac{2}{L}\operatorname{tr}(\Xi):\,\begin{array}[]{l}\xi\left(h-v_{m}\tau\right)+Z^{\intercal}y+a=-\tau\\ \xi\left(q-v_{m}1\right)-G^{\intercal}y+b=-1\\ \begin{pmatrix}\xi&(\frac{L\xi}{2}x_{0}+y)^{\intercal}\\ \frac{L\xi}{2}x_{0}+y&\Xi\end{pmatrix}\succeq 0\\ a,b\geq 0\end{array}\right\}$
Note that $\xi>0$ at any feasible solution as otherwise $\xi=0$ implies $y=0$ contradicts both of the first two constraints. Partially minimizing over $\Xi$ and dropping $a,b$ gives
$\inf_{\xi,y}\left\{\frac{2}{L}\frac{\left\lVert\frac{L\xi}{2}x_{0}+y\right\rVert^{2}}{\xi}:\,\begin{array}[]{l}\xi(h-v_{m}\tau)+Z^{\intercal}y\leq-\tau\\ \xi\left(q-v_{m}1\right)-G^{\intercal}y\leq-1\\ \xi>0\end{array}\right\}$
By Lemma 2, we know that (4) is essentially strictly feasible so that strong duality holds and that the $\inf$ can be replaced by a minimum as long as (4) is bounded.
Let $(\mu,\lambda_{\star})$ and $(\xi,y)$ denote optimal solutions to (4) and (18) respectively. Let $\Xi=\frac{1}{\xi}(\frac{L\xi}{2}x_{0}+y)(\frac{L\xi}{2}x_{0}+y)^{\intercal}$. Note that by optimality of $(\mu,\lambda_{\star})$, it must hold that $\left\langle\mu,h-v_{m}\tau-LZ^{\intercal}x_{0}\right\rangle+\left\langle\lambda_{\star},q-v_{m}1+LG^{\intercal}x_{0}\right\rangle=\frac{L}{2}\left\lVert Z\mu-G\lambda_{\star}\right\rVert^{2}$. Complementary slackness implies that
00 0 $=\left\langle\begin{pmatrix}\left\langle\mu,h-v_{m}\tau-LZ^{\intercal}x_{0}\right\rangle+\left\langle\lambda_{\star},q-v_{m}1+LG^{\intercal}x_{0}\right\rangle&Z\mu-G\lambda_{\star}\\ Z\mu-G\lambda_{\star}&\frac{2}{L}I\end{pmatrix},\begin{pmatrix}\xi&(\frac{L\xi}{2}x_{0}+y)^{\intercal}\\ \frac{L\xi}{2}x_{0}+y&\Xi\end{pmatrix}\right\rangle$
$=\left\langle\begin{pmatrix}\frac{L}{2}\left\lVert Z\mu-G\lambda_{\star}\right\rVert^{2}&Z\mu-G\lambda_{\star}\\ Z\mu-G\lambda_{\star}&\frac{2}{L}I\end{pmatrix},\begin{pmatrix}\xi&(\frac{L\xi}{2}x_{0}+y)^{\intercal}\\ \frac{L\xi}{2}x_{0}+y&\Xi\end{pmatrix}\right\rangle$
$=\frac{\xi L}{2}\left\lVert Z\mu-G\lambda_{\star}\right\rVert^{2}+\xi L\left\langle Z\mu-G\lambda_{\star},x_{0}+\frac{2y}{\xi L}\right\rangle+\frac{\xi L}{2}\left\lVert x_{0}+\frac{2y}{\xi L}\right\rVert^{2}$
$=\frac{\xi L}{2}\left\lVert Z\mu-G\lambda_{\star}+x_{0}+\frac{2y}{\xi L}\right\rVert^{2}.$
We deduce that
$\frac{-2y}{\xi L}=x_{0}+Z\mu-G\lambda_{\star}=z^{\prime}_{n}.\qed$